OpenStudy (anonymous):

if f^1(x)is the inverse of f(x) what is the value of f^-1(f(x)) a.x^-1 b.0 c.x d.1

OpenStudy (anonymous):

wait f^1 (x) is the inverse of f(x) not f^-1(x)?

OpenStudy (anonymous):

f(x) = x The inverse is switching place of f(x) and x x = f(x) since x is the same, we can say that inverse of f(x) = f(x) so f^-1 means inverse inverse of f(x) = f(x) = x answer C

OpenStudy (anonymous):

it's kinda strange when I explain it. Lemme know if I need to clarify anything

OpenStudy (anonymous):

okay I'll try that thank you

OpenStudy (anonymous):

it was right! thank you very much c:

OpenStudy (anonymous):

You're welcome :) <3